### v1

# Importance sampling for Sobol’ indices estimation

Abstract
We propose a new importance sampling framework for the estimation and analysis of Sobol’ indices. We show that a Sobol’ index defined under a reference input distribution can be consistently estimated from samples drawn from other sampling distributions by reweighting the estimator appropriately to account for the distribution change. We derive the optimal sampling distribution that minimizes the asymptotic variance and demonstrate its strong impact on estimation accuracy. Beyond variance reduction, the framework supports distributional sensitivity analysis via reverse importance sampling, enabling robust exploration of input distribution uncertainty with negligible additional computational cost.
1 Introduction
Complex computer models are now widely used across many scientific and engineering fields. Since they are often costly to evaluate, handling uncertainty and performing sensitivity analysis is particularly critical. In the context of global sensitivity analysis (GSA), where model inputs are treated as random variables, the aim is to understand, with as few model evaluations as possible, the global influence of each input or group of inputs on the variability of the output [8, 4].
To quantify this impact, a large variety of sensitivity indices has been proposed in the literature, each one associated with a different criterion. Some indices rely on higher-order moments [21, 24], others are based on divergence of measures between probability distributions [14, 18, 26], while others use contrast-based functionals or Cramér–von Mises-type distances [22, 33]. These indices allow for a refined view of the output distribution. Among all these tools, Sobol’ indices remain the most popular in practice. Originally introduced in [4], they are particularly well suited for scalar outputs, and can be generalized to vector-valued or functional outputs. Thanks to the Hoeffding decomposition [1], Sobol’ indices measure the contribution of each input to the output variance. As such, they are variance-based and only capture second-order effects.
In most practical applications, the model is a complex black-box code whose output distribution is unknown. Analytical formulas for Sobol’ indices are thus rarely available. As a result, these indices must be estimated, making their statistical estimation a key challenge in sensitivity analysis. Over the last decades, many estimation strategies have been proposed, which can be broadly classified into four categories: (i) Monte Carlo and quasi-Monte Carlo methods, including plain random sampling, low-discrepancy sequences, and nested designs [30, 20]; (ii) spectral and expansion techniques, such as the Fourier Amplitude Sensitivity Test (FAST) [3], Random Balance Designs (RBD) [13], the Effective Algorithm for Sensitivity Indices (EASI) [17], and polynomial chaos expansions [15]; (iii) the Pick–Freeze procedure, which directly estimates conditional variances via sample-based manipulations [29, 25]; and (iv) nonparametric estimators—including nearest-neighbor and kernel-based methods [11, 32, 36, 38, 35] and the rank-based estimator built on Chatterjee’s correlation coefficient [41, 38]—that require only a single input–output dataset.
The theoretical properties of methods from categories (iii) and (iv) are now well established. In particular, for first-order Sobol’ indices, consistency and asymptotic normality have been proven for kernel estimators [42, 19], the Pick–Freeze estimator [25], nearest-neighbor methods [32], and rank-based estimators [41].
Among them, the kernel estimator of [19] is efficient and achieves the asymptotic optimal variance bound for estimating the conditional second moment $\eta=\mathbb{E}[\mathbb{E}[Y\mid X]^{2}]$.However, it remains delicate to implement and computationally demanding. On the other hand, the rank-based estimator of [41] is much simpler to implement but sub-optimal and only suitable for first-order indices.
A fundamental objective in Monte-Carlo estimation is to reduce the variance of the estimator without introducing bias. One of the most widely used techniques for variance reduction is importance sampling [2] (IS), which consists of modifying the sampling distribution to obtain a more efficient estimator of the quantity of interest. The variance of the IS estimator strongly depends on the choice of the sampling distribution.
In this work, we reexamine the estimation of Sobol’ indices from the perspective of importance sampling. More specifically, we focus on the quantity $\eta$, which appears as a key component in the variance of Sobol’ estimators and is often the main term requiring accurate estimation. We introduce a novel reformulation of $\eta$ using an auxiliary probability distribution. We prove the existence of an optimal sampling distribution that can lead to zero asymptotic variance.
To the best of our knowledge, no prior work has established a theoretical connection between variance reduction through importance sampling and the estimation of Sobol’ indices. A notable exception is the contribution of [23], where the authors propose a computational approach sharing similarities with importance sampling: they introduce an auxiliary variable correlated with a given input and estimate conditional expectations using a quadrature scheme (Gauss–Hermite). Their main contribution lies in the simultaneous estimation of all first-order Sobol’ indices using only a single Monte-Carlo simulation, thereby reducing numerical effort compared to classical procedures of Sobol’ indices estimation. However, their method does not address variance minimization explicitly. In contrast, our contribution focuses on the classical aspect of importance sampling for variance reduction and provides a fully flexible probabilistic framework that allows estimation from alternative sampling distributions, and also offers computational benefits presented in Section 4.
The paper is structured as follows, Section 2 introduces Sobol’ indices and the classical importance sampling framework. Section 3 presents our main theoretical results. Section 4 illustrates these results with numerical experiments, including variance reduction analysis and a reverse importance sampling study of distributional sensitivity on a real test case. Section 5 concludes and outlines future research directions.
2 Main notation
Definition of Sobol’ indices
Consider a triplet $(Y,\mathbf{X},\mathbf{W})$ of random variables satisfying the relation $Y=f(\mathbf{X},\mathbf{W})$ for some measurable function $f$, where $Y$ is a real-valued square integrable output, $\mathbf{X}=(X_{1},\dots,X_{k})\in\mathbb{R}^{k}$ is a vector-valued input, and $\mathbf{W}$ a random vector independent from $\mathbf{X}$ with known distribution. In the context of sensitivity analysis, the so-called Sobol’ index $S_{u}$ is used to quantify the impact of a subset $u\subseteq\{1,\dots,k\}$ of the input variables on the output $Y$, and is defined as:
$S_{u}=\frac{\mathrm{Var}\left(\mathbb{E}[Y\mid\mathbf{X}_{u}]\right)}{\mathrm{Var}(Y)}=\frac{\mathbb{E}\left[\mathbb{E}[Y\mid\mathbf{X}_{u}]^{2}\right]-\mathbb{E}[Y]^{2}}{\mathrm{Var}(Y)},$ (1)
where:
  • $\mathbf{X}_{u}=(X_{j})_{j\in u}$ denotes the subvector of $\mathbf{X}$ corresponding to the index set $u$,
  • $\mathbf{X}_{\bar{u}}=(X_{j})_{j\in\bar{u}}$ is the complementary subvector, i.e., the remaining components of $\mathbf{X}$.
As explained earlier, the main challenge for inference purposes lies in handling the conditional expectation $m_{u}(\mathbf{X}_{u})=\mathbb{E}[Y\mid\mathbf{X}_{u}]$ in order to construct an estimator of
$\eta_{u}=\mathbb{E}\left[m_{u}^{2}(\mathbf{X}_{u})\right]=\mathbb{E}\left[Y\cdot m_{u}(\mathbf{X}_{u})\right]$ (2)
In a given-data setting where an i.i.d. sample $(\mathbf{X}^{(i)},Y^{(i)}),\;i=1,\dots,n$ is available, optimal estimators of $\eta_{u}$ are known to be asymptotically Gaussian, with asymptotic variance achieving the efficiency bound [19] :
$\sigma^{2}_{\mathrm{opt},u}(p)=4\mathbb{E}[m_{u}^{2}(X_{u})Y^{2}]-3\mathbb{E}\left[m_{u}^{4}(X_{u})\right]-\eta_{u}^{2}.$ (3)
As motivated in the introduction, we focus on improving the estimation of the key quantity $\eta_{u}$, which drives the variance of Sobol’ index estimators. For both theoretical and practical reasons, it is often beneficial to modify the original distribution of the input variables. In the importance sampling framework, briefly introduced in the next subsection, we consider the following setting: the input vector $\mathbf{X}$ is initially distributed according to a reference distribution $p$, under which the components are assumed to be dependent, although independence can also be assumed for analytical convenience. We then consider the use of an alternative sampling distribution $q$, potentially introducing a dependency structure.Our goal is to investigate whether it is possible to reduce the variance of the estimator of $\eta_{u}$ by judiciously choosing a sampling distribution $q$ and using samples $\mathbf{X}\sim q$ instead of $\mathbf{X}\sim p$.
To avoid any ambiguity, we explicitly indicate the distribution of the input $\mathbf{X}$ in the expectation notation. For example, $\mathbb{E}_{p}[\cdot]$ denotes the expectation when $\mathbf{X}$ is distributed according to the density $p$. The distribution $W$ is fixed and remains unchanged throughout. The distribution of the output $Y$ is only affected through changes in the distribution of the input.
Importance sampling (IS)
A classical variance-reduction technique. To estimate an expectation of the form $\mathbb{E}_{p}[f(\mathbf{X})]$, one rewrites it under a distribution $q$ as $\mathbb{E}_{q}[f(\mathbf{X})\,w(\mathbf{X})]$ where $w(\mathbf{x})=p(\mathbf{x})/q(\mathbf{x})$ is the likelihood ratio. To ensure the estimator is unbiased, the support of $q$ must include the support of $f(\mathbf{x})\,p(\mathbf{x})$. Given an i.i.d sample $\mathbf{X}^{(1)},...,\mathbf{X}^{(n)}$ drawn from $q$, the corresponding estimator is
$\hat{I}^{\mathrm{IS}}=\frac{1}{n}\sum_{i=1}^{n}f(\mathbf{X}^{(i)})\,w(\mathbf{X}^{(i)}).$ (4)
This estimator achieves zero variance if and only if $q(\mathbf{x})\propto f(\mathbf{x})\,p(\mathbf{x})$ almost-everywhere. This ideal distribution is not usable in practice because it depends on the unknown quantity to be estimated. However, adaptive procedures exist to approximate it, such as non-parametric methods [7] or parametric approaches like the cross-entropy method [34, 12].
It is worth noting that the effectiveness of IS [9] critically depends on the overlap between the support of $q$ and the regions where $p(\mathbf{x})f(\mathbf{x})$ is large. Poorly chosen auxiliary distributions can lead to high-variance or even unstable estimators due to extreme weight values. The next section presents our main contributions by extending the classical IS framework to characterize the optimal sampling distribution $q$ for estimating $\eta_{u}$ and improving Sobol’ indices estimation.
3 Main results
3.1 Input distribution change via importance sampling
We consider a modification of the input distribution, applied exclusively to the random vector $\mathbf{X}\in\mathbb{R}^{k}$. Let $p(\mathbf{x})$ denote the reference joint density of $\mathbf{X}$, for any subset of indices $u\subseteq\{1,\dots,k\}$, we let $\textbf{X}_{u}=(X_{j})_{j\in u}$ and denote by $p_{u}(\mathbf{x})$ its marginal density and introduce an auxiliary joint density $q(\mathbf{x})$, factorized as
$q(\mathbf{x})=q_{u}(\mathbf{x}_{u})\cdot q_{\bar{u}}(\mathbf{x}_{\bar{u}}\mid\mathbf{x}_{u}),$ (5)
where $q_{u}(\mathbf{x}_{u})$ and $q_{\bar{u}}(\cdot\mid\mathbf{x}_{u})$ denote the marginal and conditional densities, respectively. The corresponding importance weights are then given by:
$w(\mathbf{x})=\frac{p(\mathbf{x})}{q(\mathbf{x})}=\frac{p_{u}(\mathbf{x}_{u})\cdot p_{\bar{u}}(\mathbf{x}_{\bar{u}}\mid\mathbf{x}_{u})}{q_{u}(\mathbf{x}_{u})\cdot q_{\bar{u}}(\mathbf{x}_{\bar{u}}\mid\mathbf{x}_{u})}=w_{u}(\mathbf{x}_{u})\cdot w_{\bar{u}}(\mathbf{x}_{\bar{u}}\mid\mathbf{x}_{u}).$ (6)
Lemma
Let $q=q_{u}\cdot q_{\bar{u}}$ be any importance sampling distribution on the input $\mathbf{X}$, define
$Z_{u}:=\sqrt{w_{u}(\mathbf{X}_{u})}\cdot w_{\bar{u}}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})\cdot Y.$ (7)
Then:
$\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]=\sqrt{w_{u}(\mathbf{X}_{u})}\cdot m_{u}(\mathbf{X}_{u}).$ (8)
In particular, the following identity holds:
$\eta_{u}=\mathbb{E}_{p}\left[(\mathbb{E}_{p}[Y\mid\mathbf{X}_{u}])^{2}\right]=\mathbb{E}_{q}\left[\left(\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]\right)^{2}\right]=\mathbb{E}_{q}\left[Z_{u}\cdot\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]\right].$ (9)
Proof
See Appendix
This identity provides a reformulation of $\eta_{u}$ under the sampling distribution $q$ and can be consistently estimated using the auxiliary variable $Z_{u}$. We will show next that a suitable choice of $q$ leads to a reduced variance of the corresponding estimator.
3.2 Optimal asymptotic variance expression using importance sampling
We now investigate the case where the reference distribution of the input vector $\mathbf{X}$ can be freely modified, while the distribution of the noise variable $W$ remains fixed. We define $\phi(\mathbf{X})=\sqrt{\mathbb{E}[Y^{2}|\mathbf{X}]}$, noting that $\phi$ is model-specific and independent of the sampling distribution; its dependence on $p$ or $q$ does not need to be made explicit in the notation. Under $p$, the optimal asymptotic variance for estimating $\eta_{u}$ is given by
$\sigma^{2}_{\mathrm{opt},u}(p)=4\,\mathbb{E}_{p}\left[m_{u}^{2}(\mathbf{X}_{u})\cdot\mathbb{E}_{p}\left[\phi^{2}(\mathbf{X})\mid\mathbf{X}_{u}\right]\right]-3\,\mathbb{E}_{p}\left[m_{u}^{4}(\mathbf{X}_{u})\right]-\eta_{u}^{2}.$ (10)
We now derive the expression of the asymptotic variance under an arbitrary auxiliary distribution $q$, and compare it to the reference case.
proposition
Let $q(\mathbf{x})$ be a sampling density over $\mathbf{X}$, and let $Z_{u}$ be defined as in Lemma 1. Then the optimal asymptotic variance is given by
$\sigma^{2}_{\mathrm{opt},u}(q)=4\,\mathbb{E}_{q}\left[\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]^{2}\cdot\mathbb{E}_{q}[Z_{u}^{2}\mid\mathbf{X}_{u}]\right]-3\,\mathbb{E}_{q}\left[\left(\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]\right)^{4}\right]-\eta_{u}^{2}.$ (11)
Using Lemma 1 and substituting the expressions of $Z_{u}$, this variance can be rewritten under $p$ as
$\sigma^{2}_{\mathrm{opt},u}(q)$ $=4\,\mathbb{E}_{p}\left[w_{u}(\mathbf{X}_{u})\cdot m_{u}^{2}(\mathbf{X}_{u})\cdot\mathbb{E}_{p}\left[w_{\bar{u}}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})\cdot\phi^{2}(\mathbf{X})\mid\mathbf{X}_{u}\right]\right]$ (12)
$\quad-3\,\mathbb{E}_{p}\left[w_{u}(\mathbf{X}_{u})\cdot m_{u}^{4}(\mathbf{X}_{u})\right]-\eta_{u}^{2}.$
This expression highlights how the asymptotic variance depends explicitly on the choice of the sampling distribution $q$ through the weights $w_{u}$ and $w_{\bar{u}}$, and through the structure of $Z_{u}$. Optimizing $q$ can therefore lead to substantial variance reduction.
3.3 Optimal distribution: sequential construction
We aim to minimize the asymptotic variance $\sigma^{2}_{\mathrm{opt},u}(q)$ by optimizing the components of the sampling distribution $q$ in a sequential manner. Our strategy proceeds in two steps. First, for a fixed marginal $q_{u}$, we determine the optimal conditional density $q^{*}_{\bar{u}}(\mathbf{x}_{\bar{u}}\mid\mathbf{x}_{u})$ that minimizes the variance contribution given $\mathbf{x}_{u}$. This optimal conditional density is then substituted into the variance expression. In the second step, the resulting objective is minimized with respect to the marginal $q_{u}$. To characterize the optimal distribution, we rely on a key variational lemma recalled below.
Lemma
Let $\mathcal{D}_{k}$ denote the set of probability density functions on $\mathbb{R}^{k}$.
Then, for any $g\in\mathcal{D}_{k}$, the value $q\overset{ae}{=}g$ uniquely minimizes the functional
$q\mapsto\int_{\mathbb{R}^{k}}\frac{g^{2}(\mathbf{x})}{q(\mathbf{x})}\,d\mathbf{x}\quad,\quad q\in\mathcal{D}_{k}.$
Proof
Let $q\in\mathcal{D}_{k}$. By Jensen’s inequality applied to the convex function $t\mapsto t^{2}$, we obtain:
[…]
Where equality is achieved if and only if $q=g$ almost everywhere.
First step: optimization of $q_{\bar{u}\mid u}$
We begin by optimizing the conditional component $q_{\bar{u}\mid u}$ of $q$, assuming that $q_{u}$ is fixed. From Proposition 3.2, the conditional distribution influences the asymptotic variance $\sigma^{2}_{\mathrm{opt},u}(q)$ in eq(3.6) only through the term
$\mathbb{E}_{\begin{subarray}{c}p\end{subarray}}\left[w_{\bar{u}}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})\,\phi^{2}(\mathbf{X})\mid\mathbf{X}_{u}\right].$
Minimizing this expectation is therefore sufficient, since it is the only component of the variance that depends on $q_{\bar{u}\mid u}$. Applying Lemma 2 conditionally on $\mathbf{X}_{u}$ gives the optimal conditional density:
$q^{*}_{\bar{u}\mid u}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})=\frac{p_{\bar{u}}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})\,\phi(\mathbf{X})}{\mathbb{E}_{\begin{subarray}{c}p\end{subarray}}\left[\phi(\mathbf{X})\mid\mathbf{X}_{u}\right]},$ (13)
where the denominator ensures normalization. With this choice, the variance contribution when $q_{\bar{u}\mid u}=q^{*}_{\bar{u}\mid u}$ becomes $\left(\mathbb{E}_{p}[\phi(\mathbf{X})\mid\mathbf{X}_{u}]\right)^{2}$, instead of $\mathbb{E}_{p}[\phi^{2}(\mathbf{X})\mid\mathbf{X}_{u}]$ when $\mathbf{X}_{\bar{u}\mid u}\sim p_{\bar{u}}$ (i.e., $w_{\bar{u}\mid u}=1$). A full derivation is provided in the Appendix. Since the function $x\mapsto x^{2}$ is convex, Jensen’s inequality confirms that
$\left(\mathbb{E}_{p}[\phi(\mathbf{X})\mid\mathbf{X}_{u}]\right)^{2}\leq\mathbb{E}_{p}[\phi^{2}(\mathbf{X})\mid\mathbf{X}_{u}],$
which guarantees that the optimal conditional distribution $q^{*}_{\bar{u}\mid u}$ systematically reduces this part of the variance.
Second step: optimisation of the marginal $q_{u}$
We now fix the component $q_{\bar{u}\mid u}$, and seek to minimize the asymptotic variance over the marginal $q_{u}$. For any such fixed conditional, the sampling density is of the form $q=q_{u}\cdot q_{\bar{u}\mid u}$, and the corresponding variance reads
$\sigma_{\operatorname{opt},u}^{2}(q)=\mathbb{E}_{p}\left[w_{u}(\mathbf{X}_{u})\,S(\mathbf{X}_{u})\right]-\eta_{u}^{2},$
where the integrand $S(\mathbf{x}_{u})$ depends on the choice of $q_{\bar{u}\mid u}$. The optimal marginal minimizing this expression is given by
$q_{u}^{*}(\mathbf{x}_{u})=\frac{p(\mathbf{x}_{u})\,\sqrt{S_{q_{\bar{u}\mid u}}(\mathbf{x}_{u})}}{\mathbb{E}_{p}\left[\sqrt{S_{q_{\bar{u}\mid u}}(\mathbf{X}_{u})}\right]},\qquad\text{and}\qquad\sigma^{2}_{\operatorname{opt},u}(q)=\left(\mathbb{E}_{p}\left[\sqrt{S_{q_{\bar{u}\mid u}}(\mathbf{X}_{u})}\right]\right)^{2}-\eta_{u}^{2}.$ (14)
We distinguish below two particular cases:
  • Case A – Reference conditional: $q_{\bar{u}\mid u}=p_{\bar{u}\mid u}$
    $S_{A}(\mathbf{x}_{u})=4\,m_{u}^{2}(\mathbf{x}_{u})\,\mathbb{E}_{p}[\phi^{2}(\mathbf{X})\mid\mathbf{X}_{u}]-3\,m_{u}^{4}(\mathbf{x}_{u}).$
  • Case B – Optimal conditional: $q_{\bar{u}\mid u}=q^{*}_{\bar{u}\mid u}$
    $S_{B}(\mathbf{x}_{u})=4\,m_{u}^{2}(\mathbf{x}_{u})\,\left(\mathbb{E}_{p}[\phi(\mathbf{X})\mid\mathbf{X}_{u}]\right)^{2}-3\,m_{u}^{4}(\mathbf{x}_{u}).$
This satisfies the pointwise inequality $S_{B}(\mathbf{x}_{u})\leq S_{A}(\mathbf{x}_{u})$, shown in step one. The full chain of inequalities then reads
$\sigma^{2}_{\operatorname{opt},u}(q^{*})\leq\sigma^{2}_{\operatorname{opt},u}(q_{u}^{*}\cdot p_{\bar{u}\mid u})\leq\sigma^{2}_{\operatorname{opt},u}(p),$
this follows from Jensen’s inequality:
$\left(\mathbb{E}_{p}[\sqrt{S_{B}(\mathbf{X}_{u})}]\right)^{2}\leq\left(\mathbb{E}_{p}[\sqrt{S_{A}(\mathbf{X}_{u})}]\right)^{2}\leq\mathbb{E}_{p}[S_{A}(\mathbf{X}_{u})].$
That is, joint optimisation in Case B achieves the greatest variance reduction, demonstrating the improved efficiency obtained when both components of $q$ are optimised together.
3.4 Zero-variance optimal distribution
In the idealised case where the function $f$ is fully deterministic with all inputs controllable (i.e., there is no external noise variable $W$ and $Y=f(\mathbf{X})$), it becomes theoretically possible to achieve zero-variance in the estimation of $\eta_{u}$. This setting can be interpreted as a special case of the general framework discussed earlier. We still proceed in two steps, but expressions simplify significantly. In particular, the asymptotic variance now takes the form
$\sigma^{2}_{\mathrm{opt},u}(q)=\mathbb{E}_{p}\left[w_{u}(\mathbf{X}_{u})\left(4\,m_{u}^{2}(\mathbf{X}_{u})\cdot\mathbb{E}_{p}\left[w_{\bar{u}}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})\,Y^{2}\mid\mathbf{X}_{u}\right]-3\,m_{u}^{4}(\mathbf{X}_{u})\right)\right]-\eta_{u}^{2}.$ (15)
Applying Lemma 2 sequentially yields:
$q_{\bar{u}}^{*}(\mathbf{x}_{\bar{u}}\mid\mathbf{x}_{u})=\frac{p_{\bar{u}}(\mathbf{x}_{\bar{u}})\,f(\mathbf{x})}{m_{u}(\mathbf{x}_{u})},\qquad q_{u}^{*}(\mathbf{x}_{u})=\frac{p_{u}(\mathbf{x}_{u})\,m_{u}^{2}(\mathbf{x}_{u})}{\eta_{u}}.$ (16)
The resulting fully adapted sampling density is
$q^{*}(\mathbf{x})=q_{u}^{*}(\mathbf{x}_{u})\cdot q_{\bar{u}}^{*}(\mathbf{x}_{\bar{u}}\mid\mathbf{x}_{u})=\frac{p(\mathbf{x})\,f(\mathbf{x})\,m_{u}(\mathbf{x}_{u})}{\eta_{u}},\qquad\text{and}\qquad\sigma^{2}_{\mathrm{opt},u}(q^{*})=0.$ (17)
Hence, the density $q^{*}(\mathbf{x})$ achieves a zero-variance estimator in theory. However, this ideal result is not achievable in practice, as the optimal distribution depends on the unknown quantity $\eta_{u}$ that we aim to estimate.
4 Numerical Experiments
4.1 Analytical Illustration of Variance Reduction via IS
In this section, we illustrate the effectiveness of an IS transformation in reducing the variance $\eta_{u}$ estimators. As a benchmark example, we consider the classical Sobol’ g-function [8]:
$g(X_{1},\dots,X_{k})=\prod_{i=1}^{k}\frac{|4X_{i}-2|+a_{i}}{1+a_{i}},$ (18)
where $X_{i}\sim\mathcal{U}[0,1]$ are assumed to be independent inputs. This function is widely used in sensitivity analysis due to its non-linear, non-monotonic nature and the availability of analytical Sobol’ indices. We consider a 3-dimensional case $(k=3)$ with $a_{i}=i$, and focus on estimating the second-order Sobol’ index $S_{\{1,2\}}$, or equivalently $\eta_{\{1,2\}}$. The third input $X_{3}$ is treated as an uncontrolled variable $W$. We consider the case A in (3.3), where $\bar{u}=\emptyset$. In this setting, all quantities involved in the expression of the optimal asymptotic variance can be computed analytically, including the optimal sampling density. Specifically, the optimal density $q^{\star}$ to estimate $\eta_{12}$ is given by
$q^{*}(x_{1},x_{2})=\frac{p(x_{1},x_{2})\cdot\sqrt{S(x_{1},x_{2})}}{\mathbb{E}_{p}[\sqrt{S(X_{1},X_{2})}]},\quad\text{with}\quad S(x_{1},x_{2})=4\,m^{2}(x_{1},x_{2})\,\phi^{2}(x_{1},x_{2})-3\,m^{4}(x_{1},x_{2}).$
The conditional expectation $m(x_{1},x_{2})=\mathbb{E}[g(x_{1},x_{2},X_{3})\mid X_{1}=x_{1},X_{2}=x_{2}]$, and the squared term $\phi^{2}(x_{1},x_{2})=\mathbb{E}[g^{2}(x_{1},x_{2},X_{3})\mid X_{1}=x_{1},X_{2}=x_{2}]$, can be derived analytically as
$m(x_{1},x_{2})=\frac{(|4x_{1}-2|+1)(|4x_{2}-2|+2)}{6},\quad\phi^{2}(x_{1},x_{2})=\frac{99}{96}\cdot m^{2}(x_{1},x_{2}).$
This leads to the simplified expression
$S(x_{1},x_{2})=\frac{9}{8}\cdot m^{4}(x_{1},x_{2}),\quad\text{so that}\quad q^{*}(x_{1},x_{2})=\frac{m^{2}(x_{1},x_{2})}{\eta_{\{1,2\}}}.$
We first plot the shape of $q^{*}$ over $[0,1]^{2}$, and then track how the coefficient of variation evolves when gradually moving from $p$ to the optimal one $q^{*}$ via the interpolated family
$q_{t}(x_{1},x_{2})=(1-t)\,p(x_{1},x_{2})+t\,q^{*}(x_{1},x_{2}),\quad t\in[0,1].$
The coefficient of variation shown in Figure 1 is derived from the asymptotic variance under IS. It is defined as
$\mathrm{CV}_{\mathrm{opt}}(t)=\frac{\sigma_{\mathrm{opt},u}(q_{t})}{\eta_{\{1,2\}}},\quad\text{with}\quad\sigma^{2}_{\mathrm{opt},u}(q_{t})=\mathbb{E}_{p}\left[\frac{S(X_{1},X_{2})}{q_{t}(X_{1},X_{2})}\right]-\eta_{\{1,2\}}^{2}.$
Since all quantities are available analytically, $\mathrm{CV}_{\mathrm{opt}}(t)$ can be estimated using simple Monte-Carlo, by sampling $X\sim p$ and evaluating the integrand directly.
(a) Optimal density $q^{\star}(x_{1},x_{2})$
(b) Coefficient of variation as a function of $t$
Figure 1: Left: optimal sampling density $q^{\star}$. Right: relative dispersion of the estimator decreases as we move from $p$ to the optimal $q^{\star}$.
Parametric optimization
An alternative strategy consists of introducing a parametric family of auxiliary sampling densities for $(X_{1},X_{2})$, defined by
$q_{\alpha,\beta}(x_{1},x_{2})=\mathrm{Beta}(x_{1};\alpha_{1},\beta_{1})\cdot\mathrm{Beta}(x_{2};\alpha_{2},\beta_{2}),$
where the shape parameters $\alpha_{i},\beta_{i}$ are positive real numbers, i.e., $\alpha_{i},\beta_{i}\in\mathbb{R}_{+}^{*}$ for $i=1,2$. We restrict ourselves to the symmetric case $\alpha_{1}=\alpha_{2}=\alpha$ and $\beta_{1}=\beta_{2}=\beta$, so that the reference distribution corresponds to the uniform distribution at $(\alpha,\beta)=(1,1)$. The goal is to determine the optimal parameter pair $(\alpha^{*},\beta^{*})\in(\mathbb{R}_{+}^{*})^{2}$ that minimizes the asymptotic variance, namely
$(\alpha^{*},\beta^{*})=\arg\min_{(\alpha,\beta)\in(\mathbb{R}_{+}^{*})^{2}}\sigma^{2}_{\mathrm{opt},u}(q_{\alpha,\beta}).$
The following plot shows how the asymptotic variance changes across the Beta family and highlights the optimal region.
Figure 2: Asymptotic variance $\sigma_{\mathrm{opt}}^{2}(\alpha,\beta)$ over the Beta parameter grid (symmetric case) displayed on a logarithmic scale.
[…]
4.2 Distributional Sensitivity via Reverse Importance Sampling
We study the sensitivity of the quantity
$\eta_{u}(\theta)=\mathbb{E}_{p_{\theta}}\left[m_{u}^{2}(\mathbf{X}_{u})\right],$
with respect to changes in the input distribution $p_{\theta}$, based on a single dataset $(\mathbf{X}^{(i)},Y^{(i)})_{i=1}^{n}$ generated under $p$. This falls into the framework of reverse importance sampling; this method, originally referred to as importance sampling sensitivity analysis [6], is now often described as a reverse importance sampling mechanism [27] and exploits the fact that a single set of observations can be reused to estimate responses under multiple reference distributions. Note that this should not be confused with the reverse importance sampling estimator (or harmonic mean estimator) used in Bayesian model evidence estimation [5].
The rank estimator [41], based on the sequential correlation of outputs, takes as input a sequence $Y^{(i)}$ sorted with respect to the associated inputs $\mathbf{X}_{u}^{(i)}$, and converges to $\eta_{u}$ via the average of consecutive products. While initially defined under the distribution $p$, this estimator can be naturally extended to estimate $\eta_{u}(\theta)$ under any distribution $p_{\theta}$, using the results from Section 3.1, the convergence to $\eta_{u}(\theta)$ is guaranteed.
For each $\theta$, we define the reweighted output
$Z_{\theta,u}=\sqrt{\frac{p_{\theta,u}(\mathbf{X}_{u})}{p_{u}(\mathbf{X}_{u})}}\cdot\frac{p_{\theta,\bar{u}}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})}{p_{\bar{u}}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})}\cdot Y,$ (19)
so that the rank-based estimator becomes
$\hat{\eta}^{\text{rank}}_{u}(\theta)=\frac{1}{n-1}\sum_{i=1}^{n-1}Z_{\theta,u}^{(i)}\,Z_{\theta,u}^{(i+1)}.$ (20)
Hence, from a single dataset generated under $p$, one can efficiently explore the sensitivity of $\eta_{u}$ (and thus the associated Sobol’ index $S_{u}$) across the entire family of distributions $p_{\theta}$, without performing any additional evaluations of the black-box model.
Illustrative test case
Physical model
[…]
$w_{\boldsymbol{\theta}}(\mathbf{X})=\prod_{j=1}^{4}\frac{\operatorname{Beta}_{\theta_{j}}(X_{j}^{\mathrm{std}})}{\operatorname{Unif}[0,1](X_{j}^{\mathrm{std}})}.$ (21)
With the rank estimator in eq(4.3), the quantity $\eta_{u}(\boldsymbol{\theta})=\mathbb{E}_{p_{\boldsymbol{\theta}}}[m_{u}^{2}(\mathbf{X}_{u})]$ can be estimated for any configuration of Beta parameters $\boldsymbol{\theta}=(\theta_{1},\dots,\theta_{4})$. Two perturbation strategies are considered: (i) marginal perturbation, where only the pair $\theta_{j}=(\alpha_{j},\beta_{j})$ associated with a single input $X_{j}$ is varied while the remaining marginals remain uniform; and (ii) global perturbation, where several or all input marginals are varied simultaneously.
To keep the presentation simple, we focus on marginal perturbations. Specifically, we assess the distributional sensitivity of the contrast $\eta_{u}$ with $u=\{4\}$ by varying $\theta_{j}$ for one input variable at a time, keeping all others unchanged. This yields a family of perturbed estimates $\eta_{4}(\theta_{j})$, each revealing how the index $\eta_{4}$ responds to changes in the distribution of a specific input $X_{j}$, for $j=1,\dots,4$. All computations are based on the same reference dataset.
(a) Varying $\alpha_{1},\beta_{1}$ of $X_{1}$
(b) Varying $\alpha_{2},\beta_{2}$ of $X_{2}$
(c) Varying $\alpha_{3},\beta_{3}$ of $X_{3}$
(d) Varying $\alpha_{4},\beta_{4}$ of $X_{4}$
Figure 3: Effect of marginal distribution changes on $\eta_{4}$: each subplot shows how $\eta_{4}$ evolves as the Beta parameters of one input $X_{j}$ are varied, while others are kept uniform. All curves are computed from the same baseline sample ($N=2500$).
These results show that the sensitivity of $\eta_{4}(\theta_{j})$ strongly depends on the shape of the input distribution of $X_{j}$, with markedly different behaviors across the four inputs. Overall, these confirm that the proposed given-data IS framework offers a powerful and cost-effective mean to perform detailed distributional local sensitivity analyses of Sobol’ indices.
5 Conclusion
In this paper, we develop a new framework for the estimation of Sobol’ indices using importance sampling. We demonstrate that $\mathbb{E}\left[m_{u}^{2}(\mathbf{X}_{u})\right]$ can be estimated from any sampling distribution different from the reference one. In particular, we show that by choosing optimal densities, it is possible to achieve a lower bound on the asymptotic variance than the efficiency bound established in [19]. In some special cases, zero-variance estimation can theoretically be achieved; however, this is not feasible in practice since the optimal densities depend on unknown quantities. We also show that combining our method with the rank-based estimator [41] is particularly useful for studying the distributional sensitivity of Sobol’ indices, especially in the given-data setting where samples can be reused efficiently through the reverse importance sampling mechanism [27]. Our approach can be naturally extended to other contexts, such as the given-model setting where a computational budget is available to evaluate the black-box model. In such cases, we expect to approximate the optimal densities presented in our framework and achieve significant variance reduction. Future work may focus on constructing joint optimal densities for the simultaneous estimation of all $\eta_{u}$, rather than treating them independently.
Appendix A Lemma 1 and other derivations
Proof of Lemma 1
We compute the conditional expectation of $Z_{u}$ given $X_{u}=x_{u}$, and deduce $\eta_{u}$ by taking successive expectations under $q$.
$\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]$ $=\sqrt{w_{u}(\mathbf{X}_{u})}\cdot\mathbb{E}_{q}\left[w_{\bar{u}}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})\cdot\mathbb{E}[Y\mid\mathbf{X}]\,\middle|\,\mathbf{X}_{u}\right]$
$=\sqrt{w_{u}(\mathbf{X}_{u})}\cdot\mathbb{E}_{p}\left[\mathbb{E}[Y\mid\mathbf{X}]\,\middle|\,\mathbf{X}_{u}\right]$
$=\sqrt{w_{u}(\mathbf{X}_{u})}\cdot m_{u}(\mathbf{X}_{u}),$
and
$\mathbb{E}_{q}[\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]^{2}]=\mathbb{E}_{q}[w_{u}(\mathbf{X}_{u})m_{u}^{2}(\mathbf{X}_{u})]=\mathbb{E}_{p}[m_{u}^{2}(\mathbf{X}_{u})]$
Derivation of the optimal variance under $p_{u}$ in (3.7)
We start from the expression
$\sigma^{2}_{\mathrm{opt},u}(q)=4\,\mathbb{E}_{q}\left[\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]^{2}\cdot\mathbb{E}_{q}[Z_{u}^{2}\mid\mathbf{X}_{u}]\right]-3\,\mathbb{E}_{q}\left[\left(\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]\right)^{4}\right]-\eta_{u}^{2}.$
Using the last identity we deduce
$\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]^{2}=w_{u}(\mathbf{X}_{u})\,m_{u}^{2}(\mathbf{X}_{u}),\quad\left(\mathbb{E}_{q}[Z_{u}\mid\mathbf{X}_{u}]\right)^{4}=w_{u}^{2}(\mathbf{X}_{u})\,m_{u}^{4}(\mathbf{X}_{u}).$
For the second term $\mathbb{E}_{q_{\bar{u}}}[Z_{u}^{2}\mid\mathbf{X}_{u}]$, we write
$\mathbb{E}_{q}[Z_{u}^{2}\mid\mathbf{X}_{u}]$ $=\mathbb{E}_{q}\left[\left(\sqrt{w_{u}(\mathbf{X}_{u})}\,w_{\bar{u}}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})\,Y\right)^{2}\middle|\mathbf{X}_{u}\right]$
$=w_{u}(\mathbf{X}_{u})\cdot\mathbb{E}_{p}\left[w_{\bar{u}}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})\,\phi^{2}(\mathbf{X})\mid\mathbf{X}_{u}\right].$
Combining all terms, we obtain:
$\sigma^{2}_{\mathrm{opt},u}(q)$ $=4\,\mathbb{E}_{q}\left[w_{u}^{2}(\mathbf{X}_{u})\,m_{u}^{2}(\mathbf{X}_{u})\cdot\mathbb{E}_{p}\left[w_{\bar{u}}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})\cdot\phi^{2}(\mathbf{X})\,\middle|\,\mathbf{X}_{u}\right]\right]$
$\quad-3\,\mathbb{E}_{q}\left[w_{u}^{2}(\mathbf{X}_{u})\,m_{u}^{4}(\mathbf{X}_{u})\right]-\eta_{u}^{2}$
$=4\,\mathbb{E}_{p}\left[w_{u}(\mathbf{X}_{u})\,m_{u}^{2}(\mathbf{X}_{u})\cdot\mathbb{E}_{p}\left[w_{\bar{u}}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})\cdot\phi^{2}(\mathbf{X})\,\middle|\,\mathbf{X}_{u}\right]\right]$
$\quad-3\,\mathbb{E}_{p}\left[w_{u}(\mathbf{X}_{u})\,m_{u}^{4}(\mathbf{X}_{u})\right]-\eta_{u}^{2}.$
Computing optimal variance bounds
Plugging $q^{*}_{\bar{u}\mid u}(\mathbf{x}_{\bar{u}}\mid\mathbf{x}_{u})$, we compute the quantity that contributes to the variance $\sigma^{2}_{\mathrm{opt}}(q=(p_{u},q^{*}_{\bar{u}\mid u}))$. We obtain:
$\mathbb{E}_{p}\left[w(\mathbf{X})\cdot\phi^{2}(\mathbf{X})\mid\mathbf{X}_{u}\right]$ $=\mathbb{E}_{p}\left[\frac{p(\mathbf{X})}{q^{*}(\mathbf{X}\mid\mathbf{X}_{u})}\cdot\phi^{2}(\mathbf{X})\,\middle|\,\mathbf{X}_{u}\right]$
$=\mathbb{E}_{p}\left[\frac{\mathbb{E}_{p}[\phi(\mathbf{X})\mid\mathbf{X}_{u}]}{\phi(\mathbf{X})}\cdot\phi^{2}(\mathbf{X})\,\middle|\,\mathbf{X}_{u}\right]$
$=\mathbb{E}_{p}\left[\phi(\mathbf{X})\cdot\mathbb{E}_{p}[\phi(\mathbf{X})\mid\mathbf{X}_{u}]\,\middle|\,\mathbf{X}_{u}\right]$
$=\left(\mathbb{E}_{p}[\phi(\mathbf{X})\mid\mathbf{X}_{u}]\right)^{2}.$
Similarly for $q^{*}(\mathbf{x}_{u})$:
$\mathbb{E}_{p}\left[w(\mathbf{X}_{u})\,S(\mathbf{X}_{u})\right]-\eta_{u}^{2}$ $=\mathbb{E}_{p}\left[\frac{p(\mathbf{X}_{u})}{q^{*}(\mathbf{X}_{u})}\,S(\mathbf{X}_{u})\right]-\eta_{u}^{2}$
$=\mathbb{E}_{p}\left[\frac{\mathbb{E}_{p}[\sqrt{S(\mathbf{X}_{u})}]}{\sqrt{S(\mathbf{X}_{u})}}\,S(\mathbf{X}_{u})\right]-\eta_{u}^{2}$
$=\mathbb{E}_{p}\left[\mathbb{E}_{p}[\sqrt{S(\mathbf{X}_{u})}]\cdot\sqrt{S(\mathbf{X}_{u})}\right]-\eta_{u}^{2}$
$=\left(\mathbb{E}_{p}[\sqrt{S(\mathbf{X}_{u})}]\right)^{2}-\eta_{u}^{2}.$
Derivation of the zero-variance formula
We start from the general expression:
$\sigma^{2}_{\mathrm{opt},u}(q)=\mathbb{E}_{p}\left[w_{u}(\mathbf{X}_{u})\cdot\left(4\,m_{u}^{2}(\mathbf{X}_{u})\cdot\mathbb{E}_{p}\left[w_{\bar{u}}(\mathbf{X}_{\bar{u}}\mid\mathbf{X}_{u})\,\phi^{2}(\mathbf{X})\mid\mathbf{X}_{u}\right]-3\,m_{u}^{4}(\mathbf{X}_{u})\right)\right]-\eta_{u}^{2}.$
Now Plugging both optimal densities simultaneously:
$q_{u}^{*}(\mathbf{x}_{u})=\frac{p_{u}(\mathbf{x}_{u})\,m_{u}^{2}(\mathbf{x}_{u})}{\eta_{u}}\qquad q_{\bar{u}}^{*}(\mathbf{x}_{\bar{u}}\mid\mathbf{x}_{u})=\frac{p_{\bar{u}}(\mathbf{x}_{\bar{u}})\,f(\mathbf{x})}{m_{u}^{2}(\mathbf{x}_{u})}$
$\sigma^{2}_{\mathrm{opt},u}(q)$ $=\mathbb{E}_{p}\left[w(\mathbf{X})\cdot\left(4\,m^{2}(\mathbf{X})\cdot\mathbb{E}_{p}\left[w(\mathbf{X})\cdot\phi^{2}(\mathbf{X})\mid\mathbf{X}\right]-3\,m^{4}(\mathbf{X})\right)\right]-\eta_{u}^{2}$
$=\mathbb{E}_{p}\left[\frac{\eta_{u}}{m^{2}(\mathbf{X})}\cdot\left(4\,m^{2}(\mathbf{X})\cdot\mathbb{E}_{p}\left[\frac{m(\mathbf{X})}{f(\mathbf{X})}\cdot f^{2}(\mathbf{X})\mid\mathbf{X}\right]-3\,m^{4}(\mathbf{X})\right)\right]-\eta_{u}^{2}$
$=\mathbb{E}_{p}\left[\frac{\eta_{u}}{m^{2}(\mathbf{X})}\cdot\left(4\,m^{3}(\mathbf{X})\cdot m(\mathbf{X})-3\,m^{4}(\mathbf{X})\right)\right]-\eta_{u}^{2}$
$=\mathbb{E}_{p}\left[\frac{\eta_{u}}{m^{2}(\mathbf{X})}\cdot m^{4}(\mathbf{X})\right]-\eta_{u}^{2}$
$=\eta_{u}\cdot\mathbb{E}_{p}\left[m^{2}(\mathbf{X})\right]-\eta_{u}^{2}$
$=\eta_{u}^{2}-\eta_{u}^{2}=0.$